An object has two million more electrons than protons. What is the net charge of the object?
C??

the asked answer in C

thanks
Answer #1

net charge "Q" is the difference between all positive and negative charges in the object.

electrons-protons=2,000,000

the charge every electron inherantly possesses is e=-1.60217646 × 10-19coulombs

using this information you can say Q= e*2,000,000 for your answer.

answered by: Smith Rai
Answer #2

the object has more electrons than protons this means that the net charge is negative.
The charge a electron is -1.602×10-19C so if we have 2,000,000 of them multiply

(-1.602×10-19C)x(2,000,000)= -3.204×10-13C

this is the net charge of the object. Note is negative.

answered by: Enid